OpenStudy (anonymous):

Simplify i^12 Answers: - i -1 i 1

OpenStudy (angela210793):

=(i^2)^10=-1^10=1

myininaya (myininaya):

i^{4*3}=1 yep! gj angela

OpenStudy (angela210793):

^_^

OpenStudy (anonymous):

1